• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M.Olson

    Log entry time 08:02:41 on July 9,2007

    Entry number 208236

    keyword=Owl Shift Summary

    Owl Shift Summary:

    M.Olson, R. Michaels



    ***Production Runs*** (He3, 98uA, Q2=34fm-2)***

    4482 1M
    4483 1M
    4484 1M
    4485 1M

    ***Density Test (Current Scan, ~2 min runs)***

    4486 (95uA, good run)
    4487 (85uA, good run)
    4488 (75uA, good run)
    4489 (65uA, good run)
    4490 (55uA, good run)
    4491 (45uA, good run) * detector replay done, looks OK
    4492 (35uA, good run)
    4493 (25uA, 3 minutes, good run)
    4494 (15uA, 4 minutes, good run)
    4495 (05uA, 5 minutes, good run)

    Full t3 and coin replays done for these runs
    **********************************************

    ***Production Runs (He3, 93uA, Q2=34fm-2)***

    4496 1M ***current raised to 100uA during run, loop 1 JT opened 0.5%
    4497 1M ***current returned to 95uA
    4498 1M
    4499 1M
    4500 1M
    4501 1M ***current to 101uA during run, opened loop 1 JT by 0.5%
    4502 1M
    4503 1M ***current to 90uA, closed loop1 JT by 0.5% to lower flow
    4504 1M
    4505 1M
    4506 1M
    4507 1M
    4508 1M
    4509 1M
    4510 1M
    4511 1M
    4512 1M
    4513 1M
    4514 1M *** Beam position (x & y) fluctuating (<= 1mm) on BPM A and B
    4515 ***Beam postion fluctuating during run (y up to 5.0mm)

    Run 4515 in progress at end of shift